OpenStudy (nellynell2142):

Does anyone go to connections, who can help me with this? 1.Solve the system using elimination

jhonyy9 (jhonyy9):

???

OpenStudy (nellynell2142):

@jhonyy9 how do you add an attachment?

jhonyy9 (jhonyy9):

here below the Attach file - window - push it and using this you can attach file easy - hope this will help you

OpenStudy (nellynell2142):

3x+3y=27 x- 3y= -11 (8,1) (3,6) (6,3) (4,5)

OpenStudy (anonymous):

its 4,5for the top 3*4 = 12 3*5 = 15 and 15 + 12 =27

OpenStudy (nellynell2142):

2. Solve the system using elemination 4x-7y=3 x-7y= -15 (-12/5, 10/7) (-1,2) (6,3) (1, 1/7)

OpenStudy (nellynell2142):

Sorry it takes forever to type it I'm having a difficult time adding an attachment

jhonyy9 (jhonyy9):

@mcgrath1010 we dont give the answers just help - how you have got this result ? make calculi for more easy understanding - ok ?

jhonyy9 (jhonyy9):

Nelly - do you understand first exercise how you need solve it ?

OpenStudy (nellynell2142):

Not really

jhonyy9 (jhonyy9):

how you try it to solve - idea ?

jhonyy9 (jhonyy9):

when you need using elimination method what mean this ?

OpenStudy (nellynell2142):

You need to divide right?

jhonyy9 (jhonyy9):

3x+3y=27 x- 3y= -11 ------------ this is an easy case because than you see the terms of y assumed +3y and - 3y will result zero yes ?

jhonyy9 (jhonyy9):

Nelly - try assuming these equation term by term - ok ?

OpenStudy (nellynell2142):

Yes I see

jhonyy9 (jhonyy9):

than now just you need continue the calculi - what will get ?

OpenStudy (nellynell2142):

Ok

OpenStudy (nellynell2142):

I got b

jhonyy9 (jhonyy9):

are you here again ?

OpenStudy (nellynell2142):

Yes

jhonyy9 (jhonyy9):

so not is right b. but please write the calculi here - how you calculi - ok :?

OpenStudy (nellynell2142):

What is calculi?

jhonyy9 (jhonyy9):

how you solve this exercise ?

jhonyy9 (jhonyy9):

step by step

OpenStudy (nellynell2142):

There's too much lag on my end but thanks for the help I'll have to watch

jhonyy9 (jhonyy9):

why ? this is very easy

OpenStudy (nellynell2142):

The recordings

jhonyy9 (jhonyy9):

3x+3y=27 x- 3y= -11 ------------ assum term by term 4x 0 = 16 x = 16/4 x=4 this mean that the last choise is right ,yes ?

jhonyy9 (jhonyy9):

second system please solve it you

OpenStudy (nellynell2142):

There's a lot of lag

jhonyy9 (jhonyy9):

Nelly i like help you now and be secure that you understand how you can solve sytem same these using elimination method - ok ?

jhonyy9 (jhonyy9):

you dont like try it for i can checking your work now ?

jhonyy9 (jhonyy9):

ok. good luck bye